PLEASE GIVE ME DETAILED SOLUTION WITH ALL THE STEPS .

PLEASE  GIVE  ME  DETAILED  SOLUTION  WITH  ALL  THE  STEPS .

Grade:10

1 Answers

Arun
5 years ago
2nd part:-
(-1+a)/2, ( (-3)+b)/2
As Mid pt of AC =MID PT. OF BD 
SO, -3/2=(-1+a)/2
-6=2(-1+a)
-6=-2+2a
-4=2a
-2=a

-1=(-3+b)/2
-2=-3+b
-2+3=b
1=b
vertices of D= (-2,1)
 
 
Area of parallelogram = 2 * Area of Triangle ABC
 
hope it helps